OpenStudy (anonymous):

Please I help with this question... Which rule will find the nth term of the arithmetic sequence -58, -65, -72, -86, …?

OpenStudy (anonymous):

What's the common difference between terms?

OpenStudy (anonymous):

I think its 7 but im not sure

OpenStudy (anonymous):

That's right. If \(a_1\) is the first term of the sequence, \(a_2\) is the second, and so on, up to \(a_n\) being the \(n\)th term, you can write the sequence as \[a_1=-58\\ a_2=-58-7=-65\\ a_3=-65-7=-72\\ \vdots\\ a_n=a_{n-1}-7\] The trick now is to write \(a_{n-1}\) in terms of \(a_1\).

OpenStudy (anonymous):

\[\begin{cases}a_1=-58\\ a_2=-58-7\\ a_3=-58-7-7\\ a_4=-58-7-7-7\\ \vdots\\ a_n=-58-7-\cdots-7\end{cases}~~\Rightarrow~~\begin{cases}a_1=-58-0(7)\\ a_2=-58-1(7)\\ a_3=-58-2(7)\\ a_4=-58-3(7)\\ \vdots\\ a_n=-58-(n-1)(7)\end{cases}\] See the pattern?

OpenStudy (anonymous):

yeah I do thanks for explaining this to me

OpenStudy (anonymous):

np